OpenStudy (anonymous):

Simplify. n^6 · n^5 · n^4 ÷ n^3 · n^2 ÷ n

OpenStudy (lalaly):

n^15/n^5/n = n^7

jhonyy9 (jhonyy9):

n*6(n*5)(n*4) n*6(n*5)n*4 n*6(n*4) -----------/n = ---------/n you can simplify with n*5 and remain -------- n*3(n*2) n*5 n n*6(n*4) ------- = n*5(n*4) = n*9 n